40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13336     Entry time: Wed Sep 27 22:25:21 2017     In reply to: 13328     Reply to this: 13338
Author: gautam 
Type: Update 
Category: LSC 
Subject: DAC noise measurement (again) 

Attachment #1: Summary of results of measurements made on Friday. There is a lot in this plot, here is a breakdown:

  • I drove the excitation points of the coil output filter banks with raw time-series data downloaded during a DRMI lock with pyawg. Today during the meeting, Rana pointed out that we could just acquire median (as opposed to mean since the former is more immune to glitches during the averaging process) spectra during a lock, and then do the ifft in python to generate time series data for pyawg. Another advantage of doing it this way is that we don't need to store a large (~200MB in my case) file of 16k data for numerous channels. But since I already had this file, I decided against changing the methodology for this round of tests. Time series plots of the signals do not show any large glitches.
  • The SR785 was used in dual channel mode to acquire spectra from 2 coil driver outputs simultaneously, in the interest of saving time. Input range was set to -32dbVpk, AC coupled, which was the smallest value that worked for the given signal profile. Spectra were taken from DC-200Hz, with 801 points, 25 averages. The DB15 output of the coil driver board was connected to a DB15 breakout board, and then a BNC->Pomona mini-grabber adapter was used to connect to the SR785 input. The newly acquired linear power supplies for the GPIB box mean that spurious 60Hz harmonics were not present. 
  • Initially, I had planned to enable various bandstops from 20Hz-200Hz, to get a more complete profile of the noise. But in the end, I only used two elliptic bandstops (6th order, 60dB stopband attenuation): 60-90Hz, for which data is plotted in red and 90-200Hz, for which data is plotted in green
  • I've used the same noise model as I used here, plotted in dashed grey (summed with SR785 noise at the above input range, with input terminated via 50ohm terminator) - but had to tweak the parameters to get the curve to line up with the measurement. It looks like there is considerable variation between DAC channels, and certainly between the ITMX channels and the BS channels as groups.
  • I took the measurement in two conditions - with the coil de-whitening off (left column) and coil de-whitening on (right column). Note that the input to the excitation was acquired at the IN1 of the relevant filter bank, and since the de-whitening happens downstream of this, we don't have to do anything special.
  • In the right column, I have also plotted the LISO modelled noise, which was shown to match well with the measured curve, admittedly only for one channel (for the coil driver alone, so I am not taking into account the noise of the de-whitening board - I will fix this once I dig up that data).

Some remarks:

  1. According to this measurement, the de-whitening filters are the same on the ITMX channels and BS channels. So I don't understand the difference in the right column for BS and ITMX channels.
  2. While there is considerable variation between channels and also between ITMX and BS, there is certainly >6dB of reduction in the DAC noise when the de-whitening is engaged. However, no improvement was seen in the MICH error signal spectrum between 60-300Hz. So we have to continue to investigate other noises that can explain the noise in that band.
  3. Also, the realized improvement in DAC noise by turning on the coil de-whitening seems marginal - the low pass has gain of ~-80dB at 100Hz, but we seem to be hitting some sort of electronics noise in all channels at the level of ~100nV/rtHz (assuming the actual DAC noise doesn't degrade significantly when the digital simulated de-whitening filter is engaged).
  4. It remains to do the test for the ITMY channels.
  5. It would be useful to visualize the incoherent sum of all these channels - this is what should go into the MICH displacement NB. To be added.
  6. I'm currently loading pyawg from my user directory. Need to figure out a place to put this and add it to $PATH.

Data + code for this plot will be attached later.

Attachment 1: coilNoises.pdf  503 kB  | Hide | Hide all
coilNoises.pdf
ELOG V3.1.3-